OpenStudy (anonymous):

solve each system by substitution Please show how you got it and also how you found Y... Thanks :) -x+2y=-7 -2x-6y=-14

OpenStudy (anonymous):

Am I just doing your homework for you, or are you getting anything out of this?

OpenStudy (anonymous):

i am have some of the problems worked out it just i need to make sure i am doing it right ...but i am still doing it wrong

OpenStudy (anonymous):

the first step of substitution is to set one side equal to a variable. So- -x=-2y-7 then you plug in the new equation for the variable into the second equation 2(-2y-7)-6y=-14 (we changed the -2 to +2 because our equation is set to -x) then solve -4y-14-6y=-14 -10y=0 y=0 then you plug the new y into the first equation -x=-2(0)-7 -x=-7 x=7